# Continual Learning in Medical Imaging Analysis: A Comprehensive Review of Recent Advancements and Future Prospects

### v1

@@ 4 MR sequences: pre- and post-contrast T1w, T2w, & T2w FLAIR
RUNMC (Bloch et al. 2015), BMC (Bloch et al. 2015), HCRUDB (Lemaître et al. 2015), UCL (Litjens et al. 2014), BIDMC (Litjens et al. 2014), HK (Litjens et al. 2014), (Sivaswamy et al. 2015; Fumero et al. 2011; Orlando et al. 2020)
Zhang et al. (2023b)
CT
Abdomen multi- organ, tumor
CIS
Architectural
Swin UNETR (Hatamizadeh et al. 2021)
Supervised
Average DSC
BTCV (Landman et al. 2015), LiTS (Bilic et al. 2023), JHH (Xia et al. 2022) (private)
Ji et al. (2023)
3D CT
Whole-body organs
CIS
Architectural
nnUNet
Supervised
DSC HD95
@@
TotalSegmentator (Wasserthal et al. 2023), ChestOrgan, HNOrgan, EsoOrgan (3 private)
Bera et al. (2023)
MRI
Prostate, hippocampus, spleen
TIS, DIS
Rehearsal
Residual UNet
Supervised
DSC, average forgetting, BWT, accuracy
@@
(Prostate158 (Adams et al. 2022), NCI-ISBI (NCI ), Promise12 (Litjens et al. 2014), Decathlon (Antonelli et al. 2022)), (Drayd (Denovellis et al. 2021), HarP (Boccardi et al. 2015)), (Spleen dataset in Decathlon (NCI ))
Zhu et al. (2023)
CT, MRI
Abdomen, muscles, prostate
DIS
Regularization
U-Net
Supervised
DSC, ASSD
@@
Prostate (Liu et al. 2020; Bloch et al. 2015; Lemaître et al. 2015), abdomen (Mul ; Kavur et al. 2021), muscles (Zhu et al. 2021)
Liu et al. (2023a)
MRI
Brain tumor
CIS with domain shift
Regularization
ResNet-based 2D nnU-Net
Supervised
DSC, HD95
@@
BraTS2013 (Menze et al. 2014), TCIA (Clark et al. 2013), CBICA (Bakas et al. 2018)
Wang et al. (2023b)
Endoscopic images
Endoscopy, surgical instrument
CIS
Rehearsal + regularization
ResNet101
Supervised
mIoU
EDD2020 (Ali et al. 2021, 2020), EndoVis18 (Allan et al. 2020), EndoVis17 (Allan et al. 2019)
@@
Perkonigg et al. (2022)
CMR, CT, MRI
Cardiac segmentation, lung nodule detection, brain age estimation
TIS
Rehearsal
2D-UNet, Faster R-CNN, ResNet-50
Supervised
Dice score, AP, MAE
Cardiac (Campello et al. 2021), LIDC (Setio et al. 2017) + LNDb challenge (Pedrosa et al. 2019), IXI (IXI ) + OASIS-3 (LaMontagne et al. 2019)
@@ 7 Discussion: Towards the future
Our exploration of the CL methods for medical imaging reveals a mix of challenges and strategies. The literature reflects a struggle with traditional paradigms, notably the need for explicit task/domain identification. While some methods demand prior knowledge of domain IDs, others explore task/domain-agnostic approaches that aim to eliminate this reliance Aljundi et al. (2017); Bayasi et al. (2021). The evolving notion of task/domain boundaries adds another layer of complexity, with real-world scenarios often defying the assumption of well-separated shifts. Non-rigid transitions and blurred boundaries, as explored by González et al. (2022); Wang et al. (2023a), challenge existing CL approaches. Sample annotation, a common requirement for CL in medical imaging, introduces resource challenges, especially in histopathology. While supervised strategies dominate, Ashfahani and Pratama (2022) showcase the feasibility of unsupervised CL, using minimal labeled samples. The evaluation of CL strategies involves a diverse set of metrics, including stability (backward transfer - BWT) and plasticity (forward transfer - FWT). These metrics, encompassing aspects like average forgetting and incremental learning, facilitate nuanced comparisons. In the classification domain, applications like chest X-rays and skin disease classification employ diverse CL strategies. Latent replay mitigates catastrophic forgetting but increases resource consumption, while knowledge distillation balances stability and plasticity Li et al. (2020); Srivastava et al. (2021). Segmentation tasks, spanning brain sclerosis lesion segmentation to cardiac segmentation, rely on regularization techniques like EWC and knowledge distillation Karthik et al. (2022); Li et al. (2022). The simultaneous handling of classification and segmentation tasks, evident in studies on dermoscopic image-based skin disease classification Li et al. (2020) and abdominal organ segmentation Liu et al. (2022), introduces a trade-off between resource efficiency and overall model performance. The synthesis of literature and comparative analyses unveils the current state of CL in medical imaging. However, challenges persist, and future research should explore unsupervised CL, address labeling costs, and refine evaluation metrics. As the field evolves, understanding the dynamic interplay between medical imaging tasks and CL methodologies will be crucial for driving innovation. in Tables 13, 14 and 15, we present the overview of the selected publications in the domain of classification, segmentation, and both combined, respectively. We highlight the domain/application, dataset information, imaging modalities, continual learning scenarios, and strategies, including the learning process and methodology adopted/proposed by different studies.
Based on our critical analysis of existing literature, we discuss some of the open challenges and thus possible research directions for CL in the medical field.
@@
Availability of benchmark datasets: Contrary to other natural imaging fields, there are no standard benchmark datasets for continual learning in the medical field. Therefore, a fair comparison of different research advancements made in the field is challenging. MedMNIST is an attempt to provide an MNIST-like dataset for the medical field, however, it is limited by its non-complex nature of various sub-datasets and not mainly developed for CL. Hence, it does not offer various drift conditions that a CL technique could be explored to quantify.
@@
Intense labor requirements: Most of the continual learning approaches are supervised in nature and thus demand large annotated datasets for sequential training. This creates a bottleneck in model development especially in the histopathology domain as it is the most expensive in terms of annotation. To cope with a limited labeling budget in problems like cardiac segmentation, lung nodule detection, and brain age estimation, Perkonigg et al. (2022, 2021) explored an active learning strategy within the CL framework. The model indicates important samples and then only those samples are annotated by the experts. On the other hand, there are fully unsupervised CL approaches Ke et al. (2022); Pratama et al. (2021); Ashfahani and Pratama (2022) or self-supervised CL approaches Liu et al. (2023b) in other domains that should be explored here to tackle the annotation scarcity problem.
@@
Unexplored distributional changes: In the medical realm, drifts arising from covariance shifts are mainly explored that too by incorporating datasets from different centers or acquisition protocols for the same task. Thus, the type of drift is always sudden drift and easier to handle. On the other hand, there are many realistic unexplored drift settings. For example, drift within the center is close to a realistic scenario as it gives rise to smooth drift scenarios Perkonigg et al. (2022). Further, some drifts occur temporarily for example, Bichlmayer et al. (2022) observe one novel kind of domain shift in kidney-CAM-model data Bichlmayer et al. (2022), even if they follow a fixed acquisition protocol. The tissue structure in histopathology data of mouse kidneys shows degradation over time which causes the segmentation problem harder over time. Another very important kind of unexplored drift is “concept shift" or “concept drift" which refers to a situation where the shift causes a change in the relationship of the input variable with its target value Lu et al. (2018); Kumari and Saini (2022); Bhatt et al. (2022); Kumari (2021). For example, age group 30-50 are probable to be diabetic patients, but due to unforeseen reasons if the age group shifts to 15-30 then it is a concept drift for the diabetic patient identification task.
@@
Ethical considerations and patient privacy: Ethical considerations and patient privacy concerns are paramount in continual learning for medical imaging. The use of patient data for model training raises ethical questions regarding consent, data anonymization, and potential biases. Future research, in alignment with ethical guidelines discussed by Chee et al. (2023), should prioritize the development of frameworks that ensure responsible data usage, privacy preservation, and ethical guidelines. Striking a balance between innovation and the protection of patient rights is imperative for the sustainable progress of CL in medical imaging
Continual learning in diffusion models: The integration of continual learning into diffusion models Kazerouni et al. (2023) represents a promising avenue for advancing the capabilities of these models in dynamic and evolving environments. Diffusion models, widely employed in diverse fields, including image processing and medical imaging, often face challenges when confronted with changing data distributions. Continual learning techniques can play a pivotal role in allowing these models to adapt and accumulate knowledge over time, ensuring sustained performance in the face of evolving datasets and tasks. By incorporating continual learning, diffusion models can enhance their adaptability and generalizability, making them more resilient to variations in data characteristics and distribution shifts Gao and Liu (2023). This approach can significantly contribute to the robustness and effectiveness of diffusion models in real-world applications.
Continual learning in implicit neural representations: The exploration of continual learning within implicit neural representations Molaei et al. (2023) opens up new frontiers in leveraging the power of neural networks for dynamic learning scenarios. Implicit neural representations, commonly used in generative modeling and function approximation, can benefit immensely from continual learning techniques to adapt to new data and tasks seamlessly. As the data landscape evolves, implicit neural representations face challenges related to retaining past knowledge and efficiently incorporating new information Po et al. (2023); Yan et al. (2021b). Continual learning offers a solution to these challenges by enabling models to update their representations while preserving previously acquired knowledge. This not only enhances the model’s ability to handle changing data distributions but also supports the development of intelligent systems that can learn and evolve over time. Continual learning in implicit neural representations is instrumental for creating adaptable and intelligent models that can navigate the complexities of dynamic environments.
@@ 8 Conclusion
It has been proved and accepted by the research community that traditional machine learning models are ill-suited to handle the dynamic nature of data, and continual learning offers a promising solution. The systematic review in this manuscript provides a comprehensive overview of the state-of-the-art research in the field of continual learning in medical imaging analysis. We have explored various aspects of this evolving topic, including the challenges posed by changing data distributions, hardware, imaging protocols, data sources, tasks, and concept shifts in clinical practice and the need for models to adapt seamlessly. Through a meticulous analysis of the existing literature, we have examined the continual learning scenarios, strategies, level of supervision, experimental setup, evaluation schemes, and metrics employed to deal with the drifting nature of data in medical imaging. Furthermore, diverse applications of continual learning in medical imaging are discussed, ranging from disease classification and detection to intricate tasks of image segmentation. Each application area presents unique challenges and opportunities for research and development, and our review has shed light on the progress made in these domains. Additionally, a thorough collection and analysis of various evaluation matrices for forward and backward transfer facilitate robust evaluation and benchmarking of approaches.
In conclusion, this systematic review provides valuable insights into the current state of continual learning adaptation in medical image processing tasks, unexplored challenges, and promising future research directions. As medical imaging technology continues to advance, the development of continual learning models will be crucial for improving diagnostic accuracy, patient care, and overall healthcare outcomes.

### v3

@@ 6.3 Episode performance and others
Zhang et al.; Zhang et al. (2023d; 2021) MRI Prostate, optic cup and disc DIS Regularization U-Net Supervised DSC, ASSD RUNMC (Bloch et al. 2015), BMC (Bloch et al. 2015), HCRUDB (Lemaître et al. 2015), UCL (Litjens et al. 2014), BIDMC (Litjens et al. 2014), HK (Litjens et al. 2014), (Sivaswamy et al. 2015; Fumero et al. 2011; Orlando et al. 2020)
Zhang et al. (2023b) CT Abdomen multi- organ, tumor CIS Architectural Swin UNETR (Hatamizadeh et al. 2021) Supervised Average DSC BTCV (Landman et al. 2015), LiTS (Bilic et al. 2023), JHH (Xia et al. 2022) (private)
@@
Ji et al. (2023) 3D CT Whole-body organs CIS Architectural nnUNet Supervised DSC HD95 TotalSegmentator (Wasserthal et al. 2023), ChestOrgan, HNOrgan, EsoOrgan (3 private)
@@
Bera et al. (2023) MRI Prostate, hippocampus, spleen TIS, DIS Rehearsal Residual UNet Supervised DSC, average forgetting, BWT, accuracy (Prostate158 (Adams et al. 2022), NCI-ISBI (NCI ), Promise12 (Litjens et al. 2014), Decathlon (Antonelli et al. 2022)), (Drayd (Denovellis et al. 2021), HarP (Boccardi et al. 2015)), (Spleen dataset in Decathlon (NCI ))
@@
Zhu et al. (2023) CT, MRI Abdomen, muscles, prostate DIS Regularization U-Net Supervised DSC, ASSD Prostate (Liu et al. 2020; Bloch et al. 2015; Lemaître et al. 2015), abdomen (Mul ; Kavur et al. 2021), muscles (Zhu et al. 2021)
@@
Liu et al. (2023a) MRI Brain tumor CIS with domain shift Regularization ResNet-based 2D nnU-Net Supervised DSC, HD95 BraTS2013 (Menze et al. 2014), TCIA (Clark et al. 2013), CBICA (Bakas et al. 2018)
Wang et al. (2023b) Endoscopic images Endoscopy, surgical instrument CIS Rehearsal + regularization ResNet101 Supervised mIoU EDD2020 (Ali et al. 2021, 2020), EndoVis18 (Allan et al. 2020), EndoVis17 (Allan et al. 2019)
@@
Perkonigg et al. (2022) CMR, CT, MRI Cardiac segmentation, lung nodule detection, brain age estimation TIS Rehearsal 2D-UNet, Faster R-CNN, ResNet-50 Supervised Dice score, AP, MAE Cardiac (Campello et al. 2021), LIDC (Setio et al. 2017) + LNDb challenge (Pedrosa et al. 2019), IXI (IXI ) + OASIS-3 (LaMontagne et al. 2019)
@@ 7 Discussion: Towards the future
Our exploration of the CL methods for medical imaging reveals a mix of challenges and strategies. The literature reflects a struggle with traditional paradigms, notably the need for explicit task/domain identification. While some methods demand prior knowledge of domain IDs, others explore task/domain-agnostic approaches that aim to eliminate this reliance Aljundi et al. (2017); Bayasi et al. (2021). The evolving notion of task/domain boundaries adds another layer of complexity, with real-world scenarios often defying the assumption of well-separated shifts. Non-rigid transitions and blurred boundaries, as explored by González et al. (2022); Wang et al. (2023a), challenge existing CL approaches. Sample annotation, a common requirement for CL in medical imaging, introduces resource challenges, especially in histopathology. While supervised strategies dominate, Ashfahani and Pratama (2022) showcase the feasibility of unsupervised CL, using minimal labeled samples. The evaluation of CL strategies involves a diverse set of metrics, including stability (backward transfer - BWT) and plasticity (forward transfer - FWT). These metrics, encompassing aspects like average forgetting and incremental learning, facilitate nuanced comparisons. In the classification domain, applications like chest X-rays and skin disease classification employ diverse CL strategies. Latent replay mitigates catastrophic forgetting but increases resource consumption, while knowledge distillation balances stability and plasticity Li et al. (2020); Srivastava et al. (2021). Segmentation tasks, spanning brain sclerosis lesion segmentation to cardiac segmentation, rely on regularization techniques like EWC and knowledge distillation Karthik et al. (2022); Li et al. (2022). The simultaneous handling of classification and segmentation tasks, evident in studies on dermoscopic image-based skin disease classification Li et al. (2020) and abdominal organ segmentation Liu et al. (2022), introduces a trade-off between resource efficiency and overall model performance. The synthesis of literature and comparative analyses unveils the current state of CL in medical imaging. However, challenges persist, and future research should explore unsupervised CL, address labeling costs, and refine evaluation metrics. As the field evolves, understanding the dynamic interplay between medical imaging tasks and CL methodologies will be crucial for driving innovation. In Tables 13, 14 and 15, we present the overview of the selected publications in the domain of classification, segmentation, and both combined, respectively. We highlight the domain/application, dataset information, imaging modalities, CL scenarios, and strategies, including the learning process and methodology adopted/proposed by different studies. Based on our critical analysis of existing literature, we discuss some of the open challenges and thus possible research directions for CL in the medical field.
@@
Availability of benchmark datasets: Contrary to other natural imaging fields, there are no standard benchmark datasets for CL in the medical field. Therefore, a fair comparison of different research advancements made in the field is challenging. MedMNIST is an attempt to provide an MNIST-like dataset for the medical field, however, it is limited by its non-complex nature of various sub-datasets and not mainly developed for CL. Hence, it does not offer various drift conditions that a CL technique could be explored to evaluate and quantify.
@@
Intense labor requirements: Most of the CL approaches are supervised in nature and thus demand large annotated datasets for sequential training. This creates a bottleneck in model development especially in the histopathology domain as it is the most expensive in terms of annotation. To cope with a limited labeling budget in problems like cardiac segmentation, lung nodule detection, and brain age estimation, Perkonigg et al. (2022, 2021) explored an active learning strategy within the CL framework. The model indicates important samples and then only those samples are annotated by the experts. On the other hand, there are fully unsupervised CL approaches Ke et al. (2022); Pratama et al. (2021); Ashfahani and Pratama (2022) or self-supervised CL approaches Liu et al. (2023b) in other domains that should be explored here to tackle the annotation scarcity problem.
@@
Unexplored distributional changes: In the medical realm, drifts arising from covariance shifts are mainly explored that too by incorporating datasets from different centers or acquisition protocols for the same task. Thus, the type of drift is always “sudden drift” and easier to handle. On the other hand, there are many realistic unexplored drift settings. For example, drift within the center is close to a realistic scenario as it gives rise to smooth drift scenarios Perkonigg et al. (2022). Further, some drifts occur temporarily; for example, Bichlmayer et al. (2022) observe one novel kind of domain shift in kidney-CAM-model data Bichlmayer et al. (2022), even if they follow a fixed acquisition protocol. The tissue structure in histopathology data of mouse kidneys shows degradation over time, which causes the segmentation problem to be harder over time. Another very important kind of unexplored drift is “concept shift" or “concept drift", which refers to a situation where the shift causes a change in the relationship of the input variable with its target value Lu et al. (2018); Kumari and Saini (2022); Bhatt et al. (2022); Kumari (2021). For example, in the context of diabetic patient identification, consider a scenario where individuals aged 30-50 are deemed likely to be diabetic patients. However, if, for unforeseen reasons or based on lifestyle and food habits, there is a shift in the severity level and the age group to 20-50, it would be considered a concept drift for diabetic patient identification. This implies a fundamental change in the underlying patterns and characteristics associated with diabetic patients, thus influencing the performance and reliability of machine learning models.
@@
Ethical considerations and patient privacy: Ethical considerations and patient privacy concerns are paramount in CL for medical imaging. The use of patient data for model training raises ethical questions regarding consent, data anonymization, and potential biases. Future research, in alignment with ethical guidelines discussed by Chee et al. (2023), should prioritize the development of frameworks that ensure responsible data usage, privacy preservation, and ethical guidelines. Striking a balance between innovation and the protection of patient rights is imperative for the sustainable progress of CL in medical imaging.
Continual learning in diffusion models: The integration of CL into diffusion models Kazerouni et al. (2023) represents a promising avenue for advancing the capabilities of these models in dynamic and evolving environments. Diffusion models, widely employed in diverse fields, including image processing and medical imaging, often face challenges when confronted with changing data distributions. CL techniques can play a pivotal role in allowing these models to adapt and accumulate knowledge over time, ensuring sustained performance in the face of evolving datasets and tasks. By incorporating CL, diffusion models can enhance their adaptability and generalizability, making them more resilient to variations in data characteristics and distribution shifts Gao and Liu (2023). This approach can significantly contribute to the robustness and effectiveness of diffusion models in real-world applications.
Continual learning in implicit neural representations: The exploration of CL within implicit neural representations Molaei et al. (2023) opens up new frontiers in leveraging the power of neural networks for dynamic learning scenarios. Implicit neural representations, commonly used in generative modeling and function approximation, can benefit immensely from CL techniques to adapt to new data and tasks seamlessly. As the data landscape evolves, implicit neural representations face challenges related to retaining past knowledge and efficiently incorporating new information Po et al. (2023); Yan et al. (2021b). CL offers a solution to these challenges by enabling models to update their representations while preserving previously acquired knowledge. This not only enhances the model’s ability to handle changing data distributions but also supports the development of intelligent systems that can learn and evolve over time. CL in implicit neural representations is instrumental for creating adaptable and intelligent models that can navigate the complexities of dynamic environments.
@@ 8 Conclusion
It has been proved and accepted by the research community that traditional machine learning models are ill-suited to handle the dynamic nature of data, and CL offers a promising solution. The systematic review in this manuscript provides a comprehensive overview of the state-of-the-art research in the field of CL in medical image analysis. We have explored various aspects of this evolving topic, including the challenges posed by changing data distributions, hardware, imaging protocols, data sources, tasks, and concept shifts in clinical practice and the need for models to adapt seamlessly. Through a meticulous analysis of the existing literature, we have examined the CL scenarios, strategies, level of supervision, experimental setup, evaluation schemes, and metrics employed to deal with the drifting nature of medical image data. Furthermore, diverse applications of CL in medical image analysis are discussed, ranging from disease classification and detection to intricate tasks of image segmentation. Each application area presents unique challenges and opportunities for research and development, and our review has shed light on the progress made in these domains. Additionally, a thorough collection and analysis of various evaluation matrices for forward and backward transfer facilitate robust evaluation and benchmarking of approaches.
In conclusion, this systematic review provides valuable insights into the current state of CL adaptation in medical image processing tasks, unexplored challenges, and promising future research directions. As medical image analysis technology continues to advance, the development of CL models will be crucial for improving diagnostic accuracy, patient care, and overall healthcare outcomes.
Acknowledgments
This work was supported by the Federal Ministry of Education and Research (BMBF) of Germany under the grant no. FKZ 01IS21067A-C and the German Research Foundation (Deutsche Forschungsgemeinschaft, DFG) under the grant no. 445703531.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.